HELP please.



Jane has an interest in buying Gratise, Inc. stock. Having taken Finance 2660, she remembered to use the capital asset pricing model (CAPM) as a way of determining the asset’s risk-return relationship. She asked her stockbroker for Gratise's beta coefficient, and after some delay, she was given a number of 0.95. She then referred to some financial print media to determine return values for the DJIA (market) and the 20-year Treasury Bond as 10.0% and 2.0%,respectively. Lastly, after interviewing the CFO at Gratise, she learned that based on Gratise’s expected increase in market share and improved operation efficiency, Gratise is expected to provide a return to its shareholders of 10%, per year. Although Jane was a good student at CSU and can probably solve the problem independently, she has sought your assistance in making this investment decision. Should Jane purchase shares of Gratise stock? Explain the reason for your decision with sound financial theory.
Answer #1

Capital Asset Pricing Model Equation is     Rs = Rp + Rm-RF)*3

Where Rs is minimum required rate of return on security S

Rf = Risk free rate of return

Rm = Market Return

Beta is the systematic risk of S

In the above question:

DJIA Return is the market return Rm = 10%

Return on Treasury Bond is risk free return = 2%

Beta = 0.98

Therefore, minimum required rate of return on Gratise stcok should be = 2+(10-2)*0.98 = 2+8*0.98 = 9.84%

As per CAPM model, the stock should provide a return of at least 9.84, if it earns 9.84% or more than this it should be purchase. Shareholders are expecting a return of 10% from investment into this stock which is more than the minimum required return i.e. 9.84%.

That's why Jane's decision to invest in this stock seems to be a sound investment decision.